# Gradient descent avoids strict saddles with a simple line-search method too

### v1

@@ 3.2 The Riemannian case: general setup
Assume $x^{\ast}$ is a strict saddle point of $f$. Then, $\mathrm{Hess}f(x^{\ast})$ has at least one negative eigenvalue. Let $v$ be the unit eigenvector corresponding to such a negative eigenvalue $\lambda<0$. By Lemma 3.8, $\mathrm{D}g(x^{\ast})=I-\alpha\mathrm{Hess}f(x^{\ast})$, so $v$ is also an eigenvector of $\mathrm{D}g(x^{\ast})$ with eigenvalue $11-\alpha\lambda>$ 1. The conclusion then follows from Theorem 2.5. ∎
@@ 3.3 The Riemannian case: real-analytic retractions
In Corollary 3.9, we confirmed that the (RGD) iteration map avoids strict saddles provided it has the Luzin $N^{-1}$ property. To verify this condition, we adapt Corollary 3.1 and the associated tools from the Euclidean to the Riemannian case. We show that if the manifold and the retraction are real-analytic, then the iteration map
@@
Real-analytic maps are particularly useful in this context because, like polynomials, they do not vanish on large sets of points unless they are identically zero. We exploit this together with the fact that the map $\alpha\mapsto g_{\alpha}(x)$ is real-analytic (even if $f$ itself is only $C^{2}$) in order to conclude that, for a fixed $x$, the set of values of $\alpha$ for which $\mathrm{D}g_{\alpha}(x)$ is singular has measure zero.
@@
Let $U\subseteq{\mathbb{R}^{m}},V\subseteq{\mathbb{R}^{n}}$ be open sets. A map $F\colon U\rightarrow V$ is real-analytic if every point $x\in U$ has a neighborhood $W\subseteq U$ such that $F$ can be expressed as a convergent power series in $W$. More precisely, letting $F(x)=(F_{1}(x),\dots,F_{n}(x))$, for each coordinate function $F_{i}$ of $F$, there exists a power series centered at $x$ that converges to $F_{i}(z)$ for all $z\in W$:
@@
A map $F\colon\mathcal{M}\rightarrow\mathcal{N}$ between real-analytic manifolds is real-analytic if, for every point $x\in\mathcal{M}$, there exist real-analytic charts $(U,\varphi)$ around $x$ and $(V,\psi)$ around $F(x)$ with $F(U)\subseteq V$ such that the coordinate representation $\psi\circ F\circ\varphi^{-1}\colon\varphi(U)\to\psi(V)$ is a real-analytic map in the sense of Definition 3.10.
@@
[…]
@@
To check that the retraction is real analytic, it is often convenient to check that it admits a real-analytic extension. For example, on the unit sphere $\mathbb{S}^{n-1}=\{x\in{\mathbb{R}^{n}}:\|x\|=1\}$, the retraction $\mathrm{R}_{x}(v)=\frac{x+v}{\|x+v\|}=\frac{x+v}{\sqrt{1+\|v\|^{2}}}$ is real analytic because the map $(x,v)\mapsto\frac{x+v}{\sqrt{1+\|v\|^{2}}}$ is real analytic on ${\mathbb{R}^{n}}\times{\mathbb{R}^{n}}$ and the tangent bundle of $\mathbb{S}^{n-1}$ is a real-analytic submanifold of ${\mathbb{R}^{n}}\times{\mathbb{R}^{n}}$. Some details follow.
@@
$\{\delta\in{\mathbb{R}^{k}}\mid\mathrm{D}g_{\delta}(x)\textrm{ is singular\,}\}$
@@
$[M_{\alpha}(a)]_{ji}=e_{j}^{\top}\mathrm{D}\varphi_{\alpha}\big{(}g_{a}(x)\big{)}\big{[}\mathrm{D}g_{a}(x)[u_{i}]\big{]}.$
@@
Expanding $v=\sum_{i=1}^{n}c_{i}u_{i}$, the above is equivalent to $M_{\alpha}(a)[c]=0$ for some nonzero vector $c\in{\mathbb{R}^{n}}$.
@@
[…]
@@
Therefore, $Q_{1}\cap Q_{2}$ is empty. Then, each set is the complement of the so that they are both simultaneously open and closed. The only such subsets of ${\mathbb{R}}$ are the empty set and ${\mathbb{R}}$ itself. Since $Q_{1}$ is non-empty by assumption, it follows that $Q_{1}={\mathbb{R}}$, that is, $\mathrm{D}g_{a}(x)$ is singular for all $a\in{\mathbb{R}}$. However, this contradicts $\mathrm{D}g_{0}(x)=I$. ∎
@@ 4 Saddle avoidance with a backtracking line-search method
The results above pertain to general dynamical systems with a single, regular iteration map $g$, together with applications to GD with constant step size. In this section, we establish a saddle avoidance result for GD with a backtracking line-search method, as described in Algorithm 1. This is (Riemannian) GD with a modified version of the standard backtracking line-search method using the Armijo criterion (Nocedal and Wright, 2006, Alg. 3.1). The modification consists in forcing the selected step sizes to be non-increasing over the course of the algorithm. This effectively ensures that the step sizes stabilize (that is, are asymptotically constant) if the sequence of iterates converges. That is a key ingredient in our proof of Theorem 4.1.
Let us highlight that the gradient of $f$ need not be Lipschitz continuous. Even if it is (with some constant $L$), the algorithm is allowed to entertain (and possibly select) arbitrarily large step sizes (in particular, larger than $2/L$).
Input: $\tau,r\in(0,1),x_{0}\in\mathcal{M},\bar{\alpha}>0$
$\alpha_{0}\leftarrow\bar{\alpha}$triangleright Initialize first line-search with $\bar{\alpha}$
for $t=0,1,2,\dots$ do
         $\alpha_{t}\leftarrow\tau\alpha_{t}$
     end while
     $x_{t+1}\leftarrow\mathrm{R}_{x_{t}}(-\alpha_{t}\mathrm{grad}f(x_{t}))$
     $\alpha_{t+1}\leftarrow\alpha_{t}$triangleright Initialize next line-search with $\alpha_{t}$ rather than $\bar{\alpha}$
end for
@@
Iterate $x_{t+1}$ is obtained from $x_{t}$ by a gradient step with step size $\alpha_{t}$. By design, the sequence of step sizes is non-increasing: $\bar{\alpha}\geq\alpha_{0}\geq\alpha_{1}\geq\ldots$. Also, each $\alpha_{t}$ belongs to the discrete set $\{\bar{\alpha},\tau\bar{\alpha},\tau^{2}\bar{\alpha},\ldots\}$. Accordingly, define the iteration maps $g_{(i)}\colon\mathcal{M}\to\mathcal{M}$ for $i=0,1,2,\ldots$ corresponding to (R)GD with step sizes $\alpha_{(i)}=\tau^{i}\bar{\alpha}$, that is, $g_{(i)}(x)=\mathrm{R}_{x}(-\tau^{i}\bar{\alpha}\,\mathrm{grad}f(x))$.
@@
Assume the iterates $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 1 converge to a strict saddle point of $f$. We aim to show that $x_{0}$ belongs to a measure zero set. Let $x^{\ast}:=\lim_{t\to\infty}x_{t}$ denote the strict saddle.
The proof proceeds in two main steps. First, we show that the step sizes selected by Algorithm 1 eventually stabilize. This allows us to apply the CSMT to each iteration map $g_{(i)}$, one of which corresponds to the stabilized step size: this provides local control. Second, we use the Luzin $N^{-1}$ property of all iteration maps to enable global control.
Let us argue that $\alpha_{t}$ eventually becomes constant. The main ingredient is that the step size selected by the line-search procedure is lower bounded since the gradient of $f$ is locally Lipschitz continuous. Indeed, since $x_{t}$ converges to $x^{\ast}$, there exists some index $T\geq 0$ such that all subsequent iterates $x_{t}$ with $t\geq T$ remain in a closed ball $B_{\rho}(x^{\ast})$ of radius $\rho>0$ around $x^{\ast}$, where $\rho$ is chosen small enough so that the ball is compact. Now, consider the compact set
@@
[…]
(See for example (Boumal, 2023, Lem. 4.12); this is because any step size smaller than $\frac{2(1-r)}{L}$ satisfies the Armijo condition due to (7).) This implies that only finitely many step sizes in the interval $[\alpha_{\min},\bar{\alpha}]$ are used, and hence at least one step size is used infinitely many times. By design, there is exactly one such step size because, once a step size is discarded, it is never reconsidered.
Thus, there exist an exponent $i\geq 0$ and some index $K\geq T$ such that $x_{t+1}=g_{(i)}(x_{t})$ for all $t\geq K$. In particular,
@@ 5 RGD with constant step size: exponential retraction
[…]
@@ 5.1 Jacobi fields and the differential of the iteration map
Let $\dot{x}\in\mathrm{T}_{x}\mathcal{M}$. Let $I$ be an open interval around 00 and define a smooth curve $\sigma\colon I\to\mathcal{M}$ with $\sigma(0)=x,\sigma^{\prime}(0)=\dot{x}$. Since $g$ is $C^{1}$, we can write its differential as
@@ 5.2 The case of Hadamard manifolds
Further exploiting the favorable properties of Hadamard manifolds, we can also derive a strict saddle avoidance result for the proximal point algorithm: see Appendix B.
@@ 6 RGD with constant step size: projection retraction on spheres
[…]
@@ 7 Perspectives
  • Beyond GD and simple algorithms. Theorem 4.1 offers a first saddle avoidance result for a simple yet adaptive algorithm that includes discontinuous branching. It remains unclear how to do this for algorithms involving more complex conditional logic, mainly because of how discontinuities affect our ability to call on the CSMT. In particular, it remains open whether the standard backtracking line-search algorithm avoids saddle points.
@@ Acknowledgments
We thank Quentin Rebjock and Christopher Criscitiello for helpful discussions. Chris helped in particular with the proof of Lemma 5.9.
Appendix A Additional proofs
A.1 About the Luzin $N^{-1}$ property
@@
A.2 About measure zero sets
@@
Appendix B Proximal point algorithm on Hadamard manifolds
Recall Section 5.2 for context about saddle avoidance on Hadamard manifolds. The additional result proposed here parallels a Euclidean version stated in (Lee et al., 2019, §5.2).
Theorem B.1.
@@ Appendix B Proximal point algorithm on Hadamard manifolds
Now, let $y=g_{\mathrm{PP}}(x)$. This is the unique point such that $\mathrm{grad}q_{x}(y)=0$. Thus, $\alpha\mathrm{grad}f(y)=-\mathrm{grad}h_{x}(y)=\mathrm{Log}_{y}(x)$ and so $x=\mathrm{Exp}_{y}(\alpha\mathrm{grad}f(y))$. The right-hand side is nothing but the RGD iteration map for $-f$ at $y$, which we will denote by $g_{\mathrm{GD}}(y)$. Therefore, we have
@@
Since $\mathrm{Hess}f(x^{\ast})$ has a negative eigenvalue in the interval $[-L,0)$, we find that $\mathrm{D}g_{\mathrm{PP}}(x^{\ast})$ has an eigenvalue strictly greater than $1$, as intended. Conclude with Theorem 2.5. ∎

### v2

@@ 3.2 The Riemannian case: general setup
Assume $x^{\ast}$ is a strict saddle point of $f$. Then, $\mathrm{Hess}f(x^{\ast})$ has at least one negative eigenvalue. Let $v$ be the unit eigenvector corresponding to such a negative eigenvalue $\lambda<0$. By Lemma 3.8, $\mathrm{D}g(x^{\ast})=I-\alpha\mathrm{Hess}f(x^{\ast})$, so $v$ is also an eigenvector of $\mathrm{D}g(x^{\ast})$ with eigenvalue $1-\alpha\lambda>1$. The conclusion then follows from Theorem 2.5. ∎
@@ 3.3 The Riemannian case: real-analytic retractions
In Corollary 3.9, we confirmed that the (RGD) iteration map avoids strict saddles provided it has the Luzin $N^{-1}$ property. To verify this condition, we adapt Theorem 3.1 and the associated tools from the Euclidean to the Riemannian case. We show that if the manifold and the retraction are real analytic, then the iteration map
@@
Real-analytic maps are particularly useful in this context because, like polynomials, they do not vanish on large sets of points unless they are identically zero. We exploit this together with the fact that the map $\alpha\mapsto g_{\alpha}(x)$ is real analytic (even if $f$ itself is only $C^{2}$) in order to conclude that, for a fixed $x$, the set of values of $\alpha$ for which $\mathrm{D}g_{\alpha}(x)$ is singular has measure zero.
@@
Let $U\subseteq{\mathbb{R}^{m}},V\subseteq{\mathbb{R}^{n}}$ be open sets. A map $F\colon U\rightarrow V$ is real analytic if every point $x\in U$ has a neighborhood $W\subseteq U$ such that $F$ can be expressed as a convergent power series in $W$. More precisely, letting $F(x)=(F_{1}(x),\dots,F_{n}(x))$, for each coordinate function $F_{i}$ of $F$, there exists a power series centered at $x$ that converges to $F_{i}(z)$ for all $z\in W$:
@@
A map $F\colon\mathcal{M}\rightarrow\mathcal{N}$ between real-analytic manifolds is real analytic if, for every point $x\in\mathcal{M}$, there exist real-analytic charts $(U,\varphi)$ around $x$ and $(V,\psi)$ around $F(x)$ with $F(U)\subseteq V$ such that the coordinate representation $\psi\circ F\circ\varphi^{-1}\colon\varphi(U)\to\psi(V)$ is a real-analytic map in the sense of Definition 3.10.
@@
It is currently unclear whether the real-analyticity assumption in Theorem 3.11 is necessary. Nevertheless, some condition on the retraction is needed: without it, we can construct a counter-example. Indeed, if $f$ has a strict saddle point $x^{*}$, then one can construct a retraction that maps a large region of the tangent bundle to $x^{*}$. For instance, suppose $\mathcal{M}$ is a Hadamard manifold (i.e., complete, simply connected, with nonpositive sectional curvature). Then, the exponential retraction is globally defined and has an inverse. Choose a bounded open set $U\subseteq\mathcal{M}$ such that $\inf_{x\in U}\|\mathrm{grad}f(x)\|\geq G>0$, where $\|\cdot\|$ denotes the norm induced by the Riemannian metric. Select $0<d_{1}<d_{2}<G$. Let $q\colon[0,\infty)\to[0,1]$ be a smooth transition function such that $q(t)=1$ for $t\leq d_{1}^{2}$ and $q(t)=0$ for $t\geq d_{2}^{2}$. Then, define the retraction as
@@
To check that the retraction is real analytic, it is often convenient to check that it admits a real-analytic extension. For example, on the unit sphere $\mathbb{S}^{n-1}=\{x\in{\mathbb{R}^{n}}:\|x\|=1\}$, the retraction $\mathrm{R}_{x}(v)=\frac{x+v}{\|x+v\|}=\frac{x+v}{\sqrt{1+\|v\|^{2}}}$ is real analytic because the map $(x,v)\mapsto\frac{x+v}{\sqrt{1+\|v\|^{2}}}$ is real analytic on ${\mathbb{R}^{n}}\times{\mathbb{R}^{n}}$ and the tangent bundle of $\mathbb{S}^{n-1}$ is a real-analytic submanifold of ${\mathbb{R}^{n}}\times{\mathbb{R}^{n}}$. Some details follow (arxiv version).
@@
[…]
@@
$M_{\alpha}(a)]_{ji}=e_{j}^{\top}\mathrm{D}\varphi_{\alpha}\big(g_{a}(x)\big)\big[\mathrm{D}g_{a}(x)[u_{i}]\big].$
@@
Expanding $v=\sum_{i=1}^{n}c_{i}u_{i}$, the above is equivalent to $M_{\alpha}(a)[c]=0$ for some nonzero vector $c\in{\mathbb{R}^{n}}$. For the reverse direction, assume there exists a non-zero vector $v=\sum_{i=1}^{n}c_{i}u_{i}$ in the kernel of $M_{\alpha}(a)$. This implies that $M_{\alpha}(a)c]_{j}=0$ for all $j$. At the same time, we have that
$[M_{\alpha}(a)c]_{j}=\sum_{i=1}^{n}[M_{\alpha}(a)]_{ji}c_{i}=\sum_{i=1}^{n}c_{i}e_{j}^{\top}\mathrm{D}\varphi_{\alpha}\big(g_{a}(x)\big)\big[\mathrm{D}g_{a}(x)[u_{i}]\big]=e_{j}^{\top}\mathrm{D}\varphi_{\alpha}\big(g_{a}(x)\big)\big[\mathrm{D}g_{a}(x)[v]\big],$
hence $\mathrm{D}\varphi_{\alpha}\big(g_{a}(x)\big)\big[\mathrm{D}g_{a}(x)[v]\big]=0$, and the conclusion follows since $\mathrm{D}\varphi_{\alpha}(g_{\alpha}(x))$ is invertible.
@@
$\textrm{either }\mu\big(Z_{\alpha}\big)=0,$ $\textrm{ or }Z_{\alpha}=I_{\alpha}.$ (5)
@@
Therefore, $Q_{1}\cap Q_{2}$ is empty. Then, each set is the complement of the other one, so that they are both simultaneously open and closed. The only such subsets of $\mathbb{R}$} are the empty set and $\mathbb{R}$} itself. Since $Q_{1}$ is non-empty by assumption, it follows that $Q_{1}={\mathbb{R}}$, that is, $\mathrm{D}g_{a}(x)$ is singular for all $a\in{\mathbb{R}}$. However, this contradicts $\mathrm{D}g_{0}(x)=I$. ∎
@@ 4 Saddle avoidance with a backtracking line-search method
The results above pertain to general dynamical systems with a single, regular iteration map $g$, together with applications to GD with constant step size. In this section, we establish a saddle avoidance result for GD with a backtracking line-search method, as described in Algorithm 1. This is (Riemannian) GD with a modified version of the standard backtracking line-search method using the Armijo criterion (Nocedal and Wright, 2006, Alg. 3.1).
We emphasize that our modification only deviates from the usual line-search procedure when the gradient norm falls below an arbitrary threshold $\varepsilon>0$. In this regime, the selected step sizes become non-increasing. This effectively ensures that the step sizes stabilize (that is, are asymptotically constant) if the sequence of iterates converges. That is a key ingredient in our proof of Theorem 4.1.
Let us also highlight that the gradient of $f$ need not be Lipschitz continuous. Even if it is (with some constant $L$), the algorithm is allowed to entertain (and possibly select) arbitrarily large step sizes (in particular, larger than $2/L$).
@@
Iterate $x_{t+1}$ is obtained from $x_{t}$ by a gradient step with step size $\alpha_{t}$. Each step size $\alpha_{t}$ belongs to the discrete set $\{\bar{\alpha},\tau\bar{\alpha},\tau^{2}\bar{\alpha},\ldots\}$. Accordingly, define the iteration maps $g_{(i)}\colon\mathcal{M}\to\mathcal{M}$ for $i=0,1,2,\ldots$ corresponding to (R)GD with step sizes $\alpha_{(i)}=\tau^{i}\bar{\alpha}$, that is, $g_{(i)}(x)=\mathrm{R}_{x}(-\tau^{i}\bar{\alpha}\,\mathrm{grad}f(x))$.
@@
Assume the iterates $\{x_{t}\}_{t\geq 0}$ generated by Algorithm 1 converge to a strict saddle point of $f$. We aim to show that $x_{0}$ belongs to a measure zero set. The proof proceeds in two main steps. First, we show that the step sizes selected by Algorithm 1 eventually stabilize. This allows us to apply the CSMT to each iteration map $g_{(i)}$, one of which corresponds to the stabilized step size: this provides local control. Second, we use the Luzin $N^{-1}$ property of all iteration maps to enable global control.
Let $x^{\ast}:=\lim_{t\to\infty}x_{t}$ denote the strict saddle. Since $x_{t}$ converges to $x^{\ast}$, there exists some index $T\geq 0$ such that all subsequent iterates $x_{t}$ with $t\geq T$ remain in a closed ball $B_{\rho}(x^{\ast})$ of radius $\rho>0$ around $x^{\ast}$, where $\rho$ is chosen small enough so that: (a) the ball is compact (for example, $\rho$ strictly less than the injectivity radius of $\mathcal{M}$ at $x^{*}$), and (b) $\|\mathrm{grad}f(x)\|<\varepsilon$ for all $x\in B_{\rho}(x^{\ast})$.
We now show that for iterations $t\geq T$, the step sizes are lower-bounded, hence only a finite number is used. This follows since the gradient of $f$ is locally Lipschitz continuous. Consider the compact set
@@
$\alpha_{t}\geq\min\!\Bigg(\bar{\alpha},\frac{2\tau(1-r)}{L}\Bigg)\eqqcolon\alpha_{\min},$ $\textrm{ for all }t\geq T.$
(See for example (Boumal, 2023, Lem. 4.12); this is because any step size smaller than $\frac{2(1-r)}{L}$ satisfies the Armijo condition due to (7).) This implies that only finitely many step sizes in the interval $\alpha_{\min},\bar{\alpha}]$ are used, and hence at least one step size is used infinitely many times.
Additionally, for $t\geq T$, the step sizes are non-increasing. Indeed, since all iterates $x_{t}$ remain in $B_{\rho}(x^{\ast})$ for all $t\geq T$, and $\|\mathrm{grad}f(x)\|<\varepsilon$ for all $x\in B_{\rho}(x^{\ast})$, the mechanism in line 11 of Algorithm 1 ensures that the first step size considered when determining $\alpha_{t+1}$ is $\alpha_{t}$, rather than $\bar{\alpha}$. As a result, we have $\alpha_{t}\geq\alpha_{t+1}$ for all $t\geq T$.
Therefore, since the set of allowable step sizes is finite and the sequence is eventually non-increasing, the sequence must eventually become constant. Thus, there exist an exponent $i\geq 0$ and some index $K\geq T$ such that $x_{t+1}=g_{(i)}(x_{t})$ for all $t\geq K$. In particular,
@@ 5 RGD with constant step size: exponential retraction
$\|\mathrm{Hess}f(x)[u]\|\leq L\|u\|$ $\textrm{ for all }x\in\mathcal{M},u\in\mathrm{T}_{x}\mathcal{M}.$
@@ 5.1 Jacobi fields and the differential of the iteration map
Let $\dot{x}\in\mathrm{T}_{x}\mathcal{M}$. Let $I$ be an open interval around 0 and define a smooth curve $\sigma\colon I\to\mathcal{M}$ with $\sigma(0)=x,\sigma^{\prime}(0)=\dot{x}$. Since $g$ is $C^{1}$, we can write its differential as
@@ 5.2 The case of Hadamard manifolds
Further exploiting the favorable properties of Hadamard manifolds, we can also derive a strict saddle avoidance result for the proximal point algorithm: see Appendix C (arxiv version).
@@ 6 RGD with constant step size: projection retraction on spheres
Here, $\|\cdot\|$ denotes the Euclidean norm on $\mathcal{E}$. In general, $\mathcal{P}_{\mathcal{M}}(y)$ is a set, but it reduces to a singleton on a relevant domain so that $\mathrm{R}$ as defined earlier is indeed a retraction (possibly restricted to small $v$) (Boumal, 2023, §5.12).
@@ 7 Perspectives
  • Beyond GD and simple algorithms. Theorem 4.1 offers a first saddle avoidance result for a simple yet adaptive algorithm that includes discontinuous branching. It remains unclear how to do this for algorithms involving more complex conditional logic, mainly because of how discontinuities affect our ability to call on the CSMT. In particular, it remains open whether the standard backtracking line-search algorithm (that is, Algorithm 1 with $\varepsilon=0$) avoids saddle points.
@@ Acknowledgments
We thank Quentin Rebjock, Christopher Criscitiello and Irène Waldspurger for helpful discussions. Chris helped in particular with the proof of Lemma 5.9.
Funding and Conflicts of interests
This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00027.
The authors declare no conflicts of interest.
Appendix A About the Luzin $N^{-1}$ property
@@
Appendix B About measure zero sets
@@
Appendix C Proximal point algorithm on Hadamard manifolds
This appendix is included only in the arXiv version of this paper. Recall Section 5.2 for context about saddle avoidance on Hadamard manifolds. The additional result proposed here parallels a Euclidean version stated in (Lee et al., 2019, §5.2).
Theorem C.1.
@@ Appendix C Proximal point algorithm on Hadamard manifolds
Now, let $y=g_{\mathrm{PP}}(x)$. This is the unique point such that $\mathrm{grad}q_{x}(y)=0$. Thus, $\alpha\mathrm{grad}f(y)=-\mathrm{grad}h_{x}(y)=\mathrm{Log}_{y}(x)$ and so $x=\mathrm{Exp}_{y}(\alpha\mathrm{grad}f(y))$. The right-hand side is nothing but the RGD iteration map for f-f at $y$, which we will denote by $g_{\mathrm{GD}}(y)$. Therefore, we have
@@
Since $\mathrm{Hess}f(x^{\ast})$ has a negative eigenvalue in the interval [L,0)[-L,0), we find that $\mathrm{D}g_{\mathrm{PP}}(x^{\ast})$ has an eigenvalue strictly greater than $1$, as intended. Conclude with Theorem 2.5. ∎